OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (1, 2)? y = (x - 1)2 - 2 y = (x - 1)2 + 2 y = (x + 1)2 - 2 y = (x + 1)2 + 2

OpenStudy (anonymous):

y = (x - 1)2 + 2

OpenStudy (anonymous):

vertex = (-h,k)

OpenStudy (anonymous):

We know this because of this fact. A parabola is in a form of y = a(x-h)^2+k. The vertex is h,k. So, if plug 1,2 in, we get y = a(x-1)+2

OpenStudy (anonymous):

y = (x + 1)2 - 2 i thought it was this one because this also comes out to =2

OpenStudy (anonymous):

\[y=a(x-h)^2 +k\]

OpenStudy (anonymous):

y = (x + 1)2 - 2 i thought it was this one because this also comes out to =2

OpenStudy (anonymous):

Look at what I gave you for vertex and compare it with your equation.

OpenStudy (anonymous):

what is h here and what is k here. Does the k in the equation you provided equal to the k given in the question?

OpenStudy (anonymous):

Oh i see it should be -2!

OpenStudy (anonymous):

+2 i mean

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problem.